The Pythagorean Theorem is a2 + b2= c2. C is the side opposite the right angle. In this particular problem “C” is the unknown side.
Pythagorean Theorem in the Real-World
To get from point A to point B you must avoid walking through a pond. To avoid the pond, you must walk 34 meters south and 41 meters east. To the nearest meter, how many meters would be saved if it were possible to walk through the pond? Choose: 22, 34, 53, or 75. Show all of your work and explain how you came up with your solution.
Show all of your work.

Answer by richwmiller(17219)

34+41=75 is the distance traveled
34^2+41^2=c^2
2837=c^2
53.26=c
75-53.26 =21.74 is the distance saved
round it to the nearest meter
so the closest is 22